OpenStudy (anonymous):

Solve the equation. 0.6x + 18.65 = 22.85

OpenStudy (anonymous):

subtract 18.65 from both sides then divide both sides by 0.6

OpenStudy (anonymous):

0.6x + 18.65 = 22.85 0.6x=22.85-18.65 0.6x=4.2 x=4.2/0.6 =7

OpenStudy (anonymous):

0.6x + 18.65 equal 22.85 - 18.65 -18.65 0.6x. equal 4.20 divided by .6 x. equal. 7 answer:7